English

Six new employees, two of whom are married to each other, are to be assigned six desks that are lined up in a row. If the assignment of employees to desks is made randoml

Advertisements
Advertisements

Question

Six new employees, two of whom are married to each other, are to be assigned six desks that are lined up in a row. If the assignment of employees to desks is made randomly, what is the probability that the married couple will have nonadjacent desks?

Sum
Advertisements

Solution

Number of desk occupied by one couple = 1

Only (4 + 1) = 5 persons to be assigned.

∴ Number of ways of assigning these 5 persons = 5! × 2!

Total number of ways of assigning 6 persons = 6!

∴ Probability that a couple has adjacent desk = `(5! xx 2!)/(6!) = 1/3`

So, the probability that the married couple will have no-adjacent desks = `1 - 1/3 = 2/3`.

Hence, the required probability = `2/3`.

shaalaa.com
  Is there an error in this question or solution?
Chapter 16: Probability - Exercise [Page 296]

APPEARS IN

NCERT Exemplar Mathematics Exemplar [English] Class 11
Chapter 16 Probability
Exercise | Q 2 | Page 296

Video TutorialsVIEW ALL [1]

RELATED QUESTIONS

A coin is tossed twice, what is the probability that at least one tail occurs?


A die is thrown, find the probability of following events:

  1. A prime number will appear,
  2. A number greater than or equal to 3 will appear,
  3. A number less than or equal to one will appear,
  4. A number more than 6 will appear,
  5. A number less than 6 will appear.

A card is selected from a pack of 52 cards.

  1. How many points are there in the sample space?
  2. Calculate the probability that the card is an ace of spades.
  3. Calculate the probability that the card is
    1. an ace
    2. black card.

Three coins are tossed once. Find the probability of getting

  1. 3 heads
  2. 2 heads
  3. at least 2 heads
  4. at most 2 heads
  5. no head
  6. 3 tails
  7. exactly two tails
  8. no tail
  9. atmost two tails.

In a lottery, person chooses six different natural numbers at random from 1 to 20, and if these six numbers match with the six numbers already fixed by the lottery committee, he wins the prize. What is the probability of winning the prize in the game? [Hint: order of the numbers is not important.]


Check whether the following probabilities P(A) and P(B) are consistently defined

P(A) = 0.5, P(B) = 0.7, P(A ∩ B) = 0.6


Check whether the following probabilities P(A) and P(B) are consistently defined

P(A) = 0.5, P(B) = 0.4, P(A ∪ B) = 0.8


Fill in the blank in following table:

P(A) P(B) P(A ∩ B) P(A ∪ B)
0.35 ... 0.25 0.6

If 4-digit numbers greater than 5,000 are randomly formed from the digits 0, 1, 3, 5, and 7, what is the probability of forming a number divisible by 5 when, the digits are repeated?


A and B throw a pair of dice. If A throws 9, find B's chance of throwing a higher number.

 

Two unbiased dice are thrown. Find the probability that  neither a doublet nor a total of 8 will appear


If a letter is chosen at random from the English alphabet, find the probability that the letter is  a vowel .


If a letter is chosen at random from the English alphabet, find the probability that the letter is a consonant .


Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(ii)
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]

Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(iii) 0.7 0.06 0.05 0.04 0.03 0.2 0.1

Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(iv)
\[\frac{1}{14}\]
\[\frac{2}{14}\]
\[\frac{3}{14}\]
\[\frac{4}{14}\]
\[\frac{5}{14}\]
\[\frac{6}{14}\]
\[\frac{15}{14}\]

Two dice are thrown together. The probability that neither they show equal digits nor the sum of their digits is 9 will be


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is red or yellow and numbered 1, 2, 3 or 4


Three-digit numbers are formed using the digits 0, 2, 4, 6, 8. A number is chosen at random out of these numbers. What is the probability that this number has the same digits?


Three squares of chessboard are selected at random. The probability of getting 2 squares of one colour and other of a different colour is ______.


One mapping (function) is selected at random from all the mappings of the set A = {1, 2, 3, ..., n} into itself. The probability that the mapping selected is one to one is ______.


If the letters of the word ALGORITHM are arranged at random in a row what is the probability the letters GOR must remain together as a unit?


Four candidates A, B, C, D have applied for the assignment to coach a school cricket team. If A is twice as likely to be selected as B, and B and C are given about the same chance of being selected, while C is twice as likely to be selected as D, what are the probabilities that C will be selected?


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(A)


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(B ∩ C)


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine Probability of exactly one of the three occurs


If the letters of the word ASSASSINATION are arranged at random. Find the probability that two I’s and two N’s come together


If the letters of the word ASSASSINATION are arranged at random. Find the probability that no two A’s are coming together


Seven persons are to be seated in a row. The probability that two particular persons sit next to each other is ______.


If the probabilities for A to fail in an examination is 0.2 and that for B is 0.3, then the probability that either A or B fails is ______.


The probability that a student will pass his examination is 0.73, the probability of the student getting a compartment is 0.13, and the probability that the student will either pass or get compartment is 0.96.


The sum of probabilities of two students getting distinction in their final examinations is 1.2


C1
Probability
C2
Written Description
(a) 0.95 (i) An incorrect assignment
(b) 0.02 (ii) No chance of happening
(c) – 0.3 (iii) As much chance of happening as not
(d) 0.5 (iv) Very likely to happen
(e) 0 (v) Very little chance of happening

If 4-digit numbers greater than 5,000 are randomly formed from the digits 0, 1, 3, 5, and 7, what is the probability of forming a number divisible by 5 when, the repetition of digits is not allowed?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×